OpenStudy (anonymous):

The quotient if a number and -7 decreased by 2 is 10. Find the number.

OpenStudy (anonymous):

\[\frac{n}{-7} - 2= 10\]\[-7(\frac{n}{-7} - 2) = -7(10)\]\[n + 14 = -70\]\[n + 14 - 14 = -70 - 14\]\[n = -84\]